OpenStudy (anonymous):

2 radical 18- 3 radical 3+6 radical 3 * 2 help please

OpenStudy (anonymous):

\[2\sqrt{18}-3\sqrt{3}+6\sqrt{3^{2}}\]

OpenStudy (anonymous):

is that the question?

OpenStudy (anonymous):

2 radical 18- 3 radical 3+6 radical 3 multiplication 2 radical

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

\[(2\sqrt{18}-3\sqrt{3}+6\sqrt{3})*\sqrt{2}\]

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[(2\sqrt{9*2}-3\sqrt{3}+6\sqrt{3})*\sqrt{2}\] \[=(6\sqrt{2}+3\sqrt{3})*\sqrt{2}\] \[(6*2)+3\sqrt{6}\] \[12+3\sqrt{6}\]

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

you save me

OpenStudy (anonymous):

/3,5 *6 -(18,2 +6:3,5)/ +/-85- 38,4/

OpenStudy (anonymous):

can you help me again please??